A charter flight charges its members $400 per year. But for evry new member in excessof 60 the charge for every member is reduced by $5. What number of members leads to maximum revenue?
Answer by robertb(5830)

Let x = # of members in excess of 60.
Then for x members in excess of 60, the total number of members is 60 + x, while the unit price is 400 - 5x. The revenue is then
(60 + x)(400 - 5x), a quadratic expression. The value of x that would maximize the revenues would be the midway between the x-intercepts, or (-60 + 80)/2 = 20/2 = 10. Hence 70 members would maximize revenues.